If A and B are independent events with P(A)=1/4 and P(B)=2/3 calculate P(B|A) calculate P(A and B)

If P(AandB)=.15 and P(B)=.6 what is P(A|B)
Answer #1

the P(Bra) and Stip- Callurate P(Aund B) Now, following PCA) - ko P(B) = 20 3 =) PCA and B) = P(A) x P (B) 2 2 ol .12 PCA undP(A/B) Step-2 To (ufulete the Nowe CA and B) = 0.15 0.6 PCB) = P(A/B P(A and B) PCB) 0,15 0.6 ܠܐܝܩܝܠܩܓܝ